The angle is 64° - find the complement and the supplement
Answer by gsmani_iyer(201)


The angle = 64 degrees.
complement angle = 90-64 = 26 degrees. So 26 and 64 degrees are complement to each other. If an angle is subtracted from 90 degrees, then it is complementary angle.
Supplementary angle = 180 - 64 = 116 degrees. So 116 and 64 degrees are supplementary angles to each other. In other words, if an angle is subtracted from 180 degrees, then the difference is known as supplementary angle.
I hope this is clear to you. All the best.